OpenStudy (anonymous):

Find two positive numbers with a sum of 35. These two numbers must be the highest value possible when one is put the power of 4 and multiplied by the other to the power of 3

OpenStudy (anonymous):

Using derivatives*

OpenStudy (anonymous):

let the numbers be x and y,x>0,y>0 0<x<35,0<y<35 x+y=35 x=35-y \(let p=x^3y^4\) \(p=(35-y)^3y^4\) \[\frac{ dp }{ dy }=y^4\times3(35-y)^2(-1)+4y^3(35-y)^3=y^3(35-y)^2[-3y+4(35-y)]\] \[=y^3(35-y)^2[140-7y]\] \[y \neq0,y \neq35\] \[\frac{ dp }{ dy }=0,gives 140-7y=0,y=20\]

OpenStudy (anonymous):

when y<0,slightly \[\frac{ dp }{ dy }=(+)(+)(+)=+\] when y>0 slightly \[\frac{ dp }{ dy }=(+)(+)(-)=-\] \[\frac{ dp }{ dy }~changes~sign~as~p~ passes through y=20\] hence there is maxima at y=20 so x=35-20=15 Required numbers are 15,20

OpenStudy (anonymous):

correction when y<20, slightly and y> 20 ,slightly

OpenStudy (anonymous):

Hey, sorry for not responding earlier, but I would like to thanks you. I really appreciate your help :)

OpenStudy (anonymous):

yw